# RESEARCH ARTICLE

### v1

@@ Abstract
[ABSTRACT] This paper introduces TVB C++, a streamlined and fast C++ Back-End for The Virtual Brain (TVB), a renowned platform and a benchmark tool for full-brain simulation. TVB C++ is engineered with speed as a primary focus while retaining the flexibility and ease of use characteristic of the original TVB platform. Positioned as a complementary tool, TVB serves as a prototyping platform, whereas TVB C++ becomes indispensable when performance is paramount, particularly for large-scale simulations and leveraging advanced computation facilities like supercomputers. Developed as a TVB-compatible Back-End, TVB C++ seamlessly integrates with the original TVB implementation, facilitating effortless usage. Users can easily configure TVB C++ to execute the same code as in TVB but with enhanced performance and parallelism capabilities.
doi:
10.1093/nsr/XXXX
@@
The Virtual Brain — Whole-Brain model — C++ — Fast
@@ 1 Introduction
Despite its great versatility, TVB faces a competitive challenge in terms of computational speed when compared to other simulation platforms, especially when dealing with extensive networks or large-scale simulations. Several approaches have been proposed in the recent years to boost the performance of TVB, e.g., by porting specific models into the C-language [CITE] or interfacing with CUDA [CITE]. However, these attempts lack of the flexibility of TVB because they were restricted to specific scientific questions they wanted to address, without resulting in a general setting with the flexibility TVB has.
Here, we introduce TVB C++, a groundbreaking addition to the TVB family aimed at addressing these limitations simultaneously. Developed upon the powerful C++ programming language and the advanced Eigen numeric library [14], TVB C++ offers great computational performance without compromising on flexibility or ease of use. Unlike previous efforts —focused solely on performance—TVB C++ takes a holistic approach by retaining the same structure of the original TVB platform. This allows for TVB C++ to seamlessly integrate with TVB through Python bindings and, therefore, effortlessly complement existing workflows, offering a swift and efficient backend solution for bulk simulations where computational time is critical. With TVB C++, researchers gain access to a cutting-edge tool that empowers them to explore the intricacies of large-scale brain dynamics or conducting extensive parameter explorations, TVB C++ stands ready to help advance the field, offering unprecedented speed, performance, and versatility.
@@ 2 Background
However, the original TVB implementation —which we will refer to as TVB-Python in the following— has in computational speed its biggest drawback. To overcome slow execution speeds, two different strategies have been implemented. The first one, TVB-HPC [^2] https://wiki.ebrains.eu/bin/view/Collabs/rateml-tvb [12], automatically produces high-performance codes for CPUs and GPUs using an easy XML-based language called RateML for model specification. RateML is based on the domain-independent language LEMS [31], which allows for the declarative description of computational models using a simple XML syntax. The already existing example implementations can be easily adapted to test different models, without requiring any knowledge about algorithmic optimization. The second one, Fast_TVB (first used in [30]), is a specialized high-performance implementation of the ”Reduced Wong Wang” model [32] written in C. It makes use of several optimization strategies and a sparse memory layout to efficiently use CPU resources, which makes it possible to simulate extremely large models with millions of nodes even on a standard computer in a reasonable time. Aquilué-Llorens et al. [33] presented a CPU-based version of the TVB-AdEx model [34, 35], used to study mechanisms underlying the emergence of conscious- and unconscious-like brain state dynamics using an MPI multi-node CPU setup. More recently, [13] presented the TVB-HPC framework, a modular set of methods used here to specifically implement the TVB-AdEx model for GPU and analyze emergent dynamics, accelerating simulations and thus reducing computational resource requirements, which in turn enabled larger parameter-space explorations.
@@ 3 TVB Python
From a high-level perspective, the architecture of TVB, as illustrated in Figure 2, accommodates interaction through various interfaces tailored to specific user needs and deployment scenarios. Structurally, TVB comprises two primary components: the backend, referred to as the ”scientific computing core,” and the supporting framework, which includes the graphical user interface and data visualization tools. TVB relies on an efficient and resilient storage data system to manage both metadata information and the underlying data. Functionally, TVB backends are designed to seamlessly support various interfaces, ensuring consistent functionality regardless of the selected interface.
@@ 4 TVB C++
The architecture of TVB C++ centers on two distinct interfaces tailored to user interaction. Currently, the console interface operates independently of the storage layer, storing results solely in memory. Consequently, users utilizing the console interface must manually oversee data import and export operations. On the other hand, its back-end comprises blocks that seamlessly interface with various top application layers. The data types serve as the common language between different components such as analyzers, visualizers, simulators, and uploaders. These datatypes embody ”active data,” signifying that when TVB C++ is configured with a database, instances encapsulating data types are automatically retained.
@@ 4.8.1 Estimate BOLD signals from neural output
One option is to use convolutions with Hemodynamic Response Function (HRF) to compute the output BOLD signal, derived using generic techniques from nonlinear system identification [41, 42, 43, 44, 45]. The approach employed allows for the estimation of Volterra kernels, which elucidate the relationship between stimulus presentation and consequent hemodynamic responses. This technique represents the BOLD signal as the outcome of a convolution with a Volterra series [46], which are, essentially, Taylor expansions extended to encompass dynamical input-state-output systems and consider the impact of the input at all times in the recent past [42]. As a result, they are regarded as model-independent, capable of representing any such dynamical system. Serving as high-order extensions of linear convolution or ”smoothing,” Volterra kernels provide a nonlinear characterization of the hemodynamic response function. Consequently, they facilitate the modeling of responses to stimuli across various contexts (e.g., different rates of word presentation) and interactions among stimuli. In TVB C++ this is implemented as a special case of a Monitor class, which are tools to record significant values from the simulation, in this case using HRF kernels. As we are following TVB’s main implementation, we are also using the first-order Volterra kernel, too.
@@ 6 Summary and Discussion
PR, VJ and GD: Conceptualization, Methodology. IM, GZ, JF and GP: methodology. IM: development, software. IM, GP, GZ, GD, PR: Visualization, Investigation. PR, VJ, GP: Formal Analysis. IM, GZ, GP: Writing – original draft. GP: Funding acquisition. All authors: Validation, Data curation, Writing – review and editing
@@ Acknowledgements
IM and GP were partially funded by Grant PID2021-122136OB-C22 funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. This work has been supported by the European Union’s Horizon 2020 research and innovation program under Specific Grant Agreement No. 785907 (HBP SGA2) and Specific Grant Agreement No. 945539 (Human Brain Project SGA3).

### v2

# TVB C++: A Fast and Flexible Back-End for The Virtual Brain

@@ Abstract
This paper introduces TVB C++, a streamlined and fast C++ Back-End for The Virtual Brain (TVB), a renowned platform and a benchmark tool for full-brain simulation. TVB C++ is engineered with speed as a primary focus while retaining the flexibility and ease of use characteristic of the original TVB platform. Positioned as a complementary tool, TVB serves as a prototyping platform, whereas TVB C++ becomes indispensable when performance is paramount, particularly for large-scale simulations and leveraging advanced computation facilities like supercomputers. Developed as a TVB-compatible Back-End, TVB C++ seamlessly integrates with the original TVB implementation, facilitating effortless usage. Users can easily configure TVB C++ to execute the same code as in TVB but with enhanced performance and parallelism capabilities.
@@
The Virtual Brain, Whole-Brain model, C++, Fast
@@ 1 Introduction
Despite its great versatility, TVB faces a competitive challenge in terms of computational speed when compared to other simulation platforms, especially when dealing with extensive networks or large-scale simulations. Several approaches have been proposed in recent years to boost the performance of TVB, e.g., by porting specific models into the C-language [CITE] or interfacing with CUDA [CITE]. However, these attempts lacked the flexibility of TVB because they were restricted to specific scientific questions they wanted to address, without resulting in a general setting with the flexibility TVB has.
Here, we introduce TVB C++, a groundbreaking addition to the TVB family aimed at addressing these limitations simultaneously. Developed upon the powerful C++ programming language and the advanced Eigen numeric library [14], TVB C++ offers great computational performance without compromising on flexibility or ease of use. Unlike previous efforts —focused solely on performance—TVB C++ takes a holistic approach by retaining the same structure of the original TVB platform. This allows for TVB C++ to seamlessly integrate with TVB through Python bindings and, therefore, effortlessly complement existing workflows, offering a swift and efficient backend solution for bulk simulations where computational time is critical. With TVB C++, researchers gain access to a cutting-edge tool that empowers them to explore the intricacies of large-scale brain dynamics or conduct extensive parameter explorations, TVB C++ stands ready to help advance the field, offering unprecedented speed, performance, and versatility.
@@ 2 Background
However, the original TVB implementation —which we will refer to as TVB-Python in the following— has in computational speed its biggest drawback. To overcome slow execution speeds, two different strategies have been implemented. The first one, TVB-HPC [^2] https://wiki.ebrains.eu/bin/view/Collabs/rateml-tvb [12], automatically produces high-performance codes for CPUs and GPUs using an easy XML-based language called RateML for model specification. RateML is based on the domain-independent language LEMS [31], which allows for the declarative description of computational models using a simple XML syntax. The already existing example implementations can be easily adapted to test different models, without requiring any knowledge about algorithmic optimization. The second one, Fast_TVB (first used in [30]), is a specialized high-performance implementation of the "Reduced Wong Wang" model [32] written in C. It makes use of several optimization strategies and a sparse memory layout to efficiently use CPU resources, which makes it possible to simulate extremely large models with millions of nodes even on a standard computer in a reasonable time. Aquilué-Llorens et al. [33] presented a CPU-based version of the TVB-AdEx model [34, 35], used to study mechanisms underlying the emergence of conscious- and unconscious-like brain state dynamics using an MPI multi-node CPU setup. More recently, [13] presented the TVB-HPC framework, a modular set of methods used here to specifically implement the TVB-AdEx model for GPU and analyze emergent dynamics, accelerating simulations and thus reducing computational resource requirements, which in turn enabled larger parameter-space explorations.
@@ 3 TVB Python
From a high-level perspective, the architecture of TVB, as illustrated in Figure 2, accommodates interaction through various interfaces tailored to specific user needs and deployment scenarios. Structurally, TVB comprises two primary components: the backend, referred to as the "scientific computing core," and the supporting framework, which includes the graphical user interface and data visualization tools. TVB relies on an efficient and resilient storage data system to manage both metadata information and the underlying data. Functionally, TVB backends are designed to seamlessly support various interfaces, ensuring consistent functionality regardless of the selected interface.
@@ 4 TVB C++
The architecture of TVB C++ centers on two distinct interfaces tailored to user interaction. Currently, the console interface operates independently of the storage layer, storing results solely in memory. Consequently, users utilizing the console interface must manually oversee data import and export operations. On the other hand, its back-end comprises blocks that seamlessly interface with various top application layers. The data types serve as the common language between different components such as analyzers, visualizers, simulators, and uploaders. These datatypes embody "active data," signifying that when TVB C++ is configured with a database, instances encapsulating data types are automatically retained.
@@ 4.8.1 Estimate BOLD signals from neural output
One option is to use convolutions with Hemodynamic Response Function (HRF) to compute the output BOLD signal, derived using generic techniques from nonlinear system identification [41, 42, 43, 44, 45]. The approach employed allows for the estimation of Volterra kernels, which elucidate the relationship between stimulus presentation and consequent hemodynamic responses. This technique represents the BOLD signal as the outcome of a convolution with a Volterra series [46], which are, essentially, Taylor expansions extended to encompass dynamical input-state-output systems and consider the impact of the input at all times in the recent past [42]. As a result, they are regarded as model-independent, capable of representing any such dynamical system. Serving as high-order extensions of linear convolution or "smoothing," Volterra kernels provide a nonlinear characterization of the hemodynamic response function. Consequently, they facilitate the modeling of responses to stimuli across various contexts (e.g., different rates of word presentation) and interactions among stimuli. In TVB C++ this is implemented as a special case of a Monitor class, which are tools to record significant values from the simulation, in this case using HRF kernels. As we are following TVB’s main implementation, we are also using the first-order Volterra kernel, too.
@@ 6 Summary and Discussion
PR, VJ and GD: conceptualization, methodology. IM, GZ, JF, MS and GP: methodology. IM: development, software. IM, GP, GZ, GD, MS, PR: Visualization, Investigation. PR, VJ, GP: Formal Analysis. IM, GZ, GP: Writing – original draft. GP: Funding acquisition. All authors: Validation, Data curation, Writing – review and editing
@@ Acknowledgements
IM and GP were partially funded by Grant PID2021-122136OB-C22 funded by MCIN/AEI/10.13039/501100011033 and by ERDF A way of making Europe. This work has been supported by the European Union’s Horizon 2020 research and innovation program under Specific Grant Agreement No. 785907 (HBP SGA2) and Specific Grant Agreement No. 945539 (Human Brain Project SGA3). This work was supported by the Virtual Research Environment at the Charité Berlin – a node of EBRAINS Health Data Cloud. Part of computation has been performed on the HPC for Research cluster of the Berlin Institute of Health. PR acknowledges support by EU Horizon Europe program Horizon EBRAINS2.0 (101147319), Virtual Brain Twin (101137289), EBRAINS-PREP 101079717, AISN – 101057655, EBRAIN-Health 101058516, Digital Europe TEF-Health 101100700; German Research Foundation SFB 1436 (project ID 425899996); SFB 1315 (project ID 327654276); SFB 936 (project ID 178316478; SFB-TRR 295 (project ID 424778381); SPP Computational Connectomics RI 2073/6-1, RI 2073/10-2,RI 2073/9-1; DFG Clinical Research Group BECAUSE-Y 504745852, PHRASE Horizon EIC grant; 101058240; Berlin Institute of Health & Foundation Charité,